The boy throws the ball vertically, with speed v0 = 4m / s. The mass of the ball is m = 800 g. a) What impulse is communicated to the boy and the cart while throwing the ball? D) Why does the boy and the cart not move? Who hinders them? c) Knowing that the interaction of the boy with the ball during the throw lasts delta t = 0.045 s, calculate the additional compressive force exerted on the ground during this time.


1
Expert's answer
2021-01-29T05:20:26-0500

a)


"p=mv=0.8\cdot4=3.2\ (kg\cdot m/s)" . Answer


b)


The surface of the Earth hinders movement.


c)


"F=\frac{\Delta p}{\Delta t}=\frac{3.2}{0.045}=71\ (N)" . Answer